A certain shade of blue paint is made by mixing 11

quarts of blue paint with 5 quarts of white paint. If you need a total of 16.25 gallons of this shade of blue paint, how much of each color should you mix?


1
Expert's answer
2021-03-31T09:27:13-0400

1 gallon = 4 Quartz

11 Quartzs=2.75 gallons

4 Quartzs=1.25 gallons

Let quantity of blue paint=x

Let quantity of white paint=Y

Ratio of X to Y = 2.75 : 1.25

If 16.25 gallons are needed

Blue = (2.75/4)*16.25=11.171875 gallons

= 44.6875 Quartz


White = (1.25/4)*16.25=5.078125 gallons

=20.3125Quartz